OpenStudy (anonymous):

Phil has 11 stamps of denominations $0.37 and $0.23. If the total value of the stamps is $3.23 how many $0.23 stamps does Phil have?

OpenStudy (anonymous):

i would use a system of equations. x represents # of 23 cent stamps, and y 37 cent. You also know the total is 11. So: x+y=11 .23*(x)+.37*(y)=3.23 Solve this system of equations.

OpenStudy (anonymous):

Is this answer correct? x=14.0435-1.6087y

OpenStudy (anonymous):

sry, i dont think so. I got 6 23 cent stamps, and 5 37 cent stamps

OpenStudy (anonymous):

>x+y=11 >>x=11-y Replace all x in the other equation with (11-y): >.23*(x)+.37*(y)=3.23 >>.23*(11-y)+.37*(y)=3.23 and so on..

OpenStudy (anonymous):

So is it 5?